OpenStudy (anonymous):

HELP!!! i need to finish this lesson! Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

Y= mx+b m=slope b=y-intercept Ok, you know your y-intercept right of the bat. do you know what that is? You also will find your slope through the two coordinates.

OpenStudy (anonymous):

Ok, and then what?

OpenStudy (anonymous):

@mramirez686

OpenStudy (anonymous):

Well you know what your y-intercept is right?

OpenStudy (anonymous):

no.. @mramirez686

OpenStudy (anonymous):

y-intercept is when the line crosses the y axis. This is when x = 0. That coordinate is given.

OpenStudy (anonymous):

ok i get that

OpenStudy (anonymous):

to find your slope, the equation is \[\frac{y_{1}-y_{2} }{x _{1}-x _{2} }= m\]

OpenStudy (anonymous):

So, right now, you know y=mx-3. next step is to get slope.

OpenStudy (anonymous):

(3,6) \[\frac{ 6-(-3) }{ 3-0 }=\frac{ 9 }{ 3 }\] (0,-3)

OpenStudy (anonymous):

do you understand how I got that?

OpenStudy (anonymous):

So your slope is 9/3 or simply 3. Now you have your equation: y=3x-3

OpenStudy (anonymous):

To check if you are right, insert a given coordinate. In other words, replace the x and y. (3,6) y=3x-3 (6)=3(3)-3 6=9-3 6=6 ...and you are good. Any questions?